# A Neuroscience-Inspired Dual-Process Model of Compositional Generalization

### v1

@@ Abstract
[…]
@@ 1 Introduction
Humans can systematically generalize to novel situations by composing familiar concepts in new ways Lake et al. (2017); Spelke and Kinzler (2007); Fodor (1975). This compositional flexibility is most evident in natural language, but is present more broadly in abstract reasoning and even everyday thought, beginning early in life Pomiechowska et al. (2024); Piantadosi et al. (2018); Dautriche and Chemla (2025). As a concrete example, knowing how to “jump”, “turn left”, and do something “twice” allows us to understand “jump twice” or “turn left after jumping.” This capacity for compositional generalization is a hallmark of human intelligence but has proven difficult to replicate in artificial systems Fodor and Pylyshyn (1988); Lake and Baroni (2018). While large language models (LLMs) have achieved remarkable success on various NLP tasks, their ability to systematically generalize, especially on tasks requiring strong compositional reasoning, remains limited Xu et al. (2024); Wang et al. (2024).
Survival in complex environments requires precisely these complementary cognitive abilities: the capacity to rapidly recognize familiar patterns and the ability to flexibly recombine known concepts into novel configurations. The mammalian brain achieves this balance through specialized neural systems that have evolved to handle different aspects of information processing. The hippocampus rapidly encodes episodic experiences, while the neocortex gradually extracts statistical regularities across many encounters, forming what neuroscientists call complementary learning systems (CLS) McClelland et al. (1995); Kumaran et al. (2016). This division enables both quick adaptation to new situations and stable accumulation of generalizable knowledge.
During cognitive processing, these systems engage in a nuanced interplay. The neocortex, basal ganglia, and related structures support intuitive, automatic processing (System 1), while a specialized circuit between the hippocampus (HPC) and prefrontal cortex (PFC) enables deliberate, structured reasoning (System 2) (Figure 1). The HPC-PFC complex is particularly crucial for compositional thinking. The hippocampus rapidly binds episodic elements to form cognitive maps of specific environments Moscovitch et al. (2016); Sun et al. (2025); George et al. (2021); Whittington et al. (2020), while the PFC, through iterative interactions with these episodic representations, extracts regularities across experiences to generate abstract schemas Bein and Niv (2025). This complementary process allows the HPC-PFC circuit to orchestrate systematic application of learned knowledge through step-by-step reasoning Koechlin et al. (2003); Tse et al. (2007); Euston et al. (2012); Preston and Eichenbaum (2013). These schemas, consolidated abstractions derived from experience, provide reusable building blocks that can be recombined to solve novel problems.
Figure 1: Dual-process systems in the brain. Left: Brain regions associated with different cognitive processing systems. The hippocampus (HPC) and prefrontal cortex (PFC) form a complex supporting deliberate reasoning (System 2), while the wider neocortex and related regions support intuitive processing (System 1). Right: Functional characteristics of each system. System 1 (Neocortex & related brain regions) is characterized by fast, intuitive pattern recognition and statistical learning. System 2 (HPC-PFC complex) enables slow, deliberate structured reasoning and schema extraction. These complementary neural systems provide the biological foundation for compositional generalization in human cognition.
Drawing on these neuroscientific insights, we present Mirage (Meta-Inference with Rules and Abstractions from Generalized Experience), a dual-process computational framework designed to achieve robust compositional generalization. Mirage mirrors the brain’s functional organization through two complementary systems:
@@
    System 1: Meta-Transformer-based Decomposer: A standard Transformer architecture Vaswani et al. (2017) that performs fast, parallel pattern recognition analogous to neocortical processing.
@@
    System 2: HPC-PFC-inspired Schema Engine: A deliberate reasoning component that extracts reusable schemas, assigns priorities to resolve ambiguities, creates temporary bindings for schema arguments, and executes structured transformations, mirroring HPC-PFC functionality.
These systems operate in coordinated iteration: System 2 identifies relevant schemas and their priorities, injects this structured information into System 1’s context window, and the Transformer uses this schema-augmented input to decompose complex problems into manageable subproblems. After each processing step by System 1, System 2 manages schema application and placeholder substitution, refining the representation for subsequent Transformer processing. This iterative schema-guided refinement allows Mirage to systematically tackle compositional tasks that would otherwise remain intractable.
We evaluate Mirage on the SCAN benchmark Lake and Baroni (2018), which is specifically designed to test compositional generalization. Our key contributions are: (1) The Mirage framework, which integrates a fast, intuitive Transformer-based processor with a deliberate, rule-based schema mechanism in a neuroscience-inspired architecture; (2) Demonstration of strong compositional generalization on SCAN, achieving $>$99% accuracy across all splits; (3) Ablation studies confirming that Mirage’s compositional capabilities critically depend on schema priority management, iterative refinement, and the quality of extracted schemas.
Figure 2: Mirage architecture showing the interplay between its two complementary systems inspired by neuroscience: (A) System 1 (Neural Decomposer) trained on diverse artificial grammars implements neocortex-like pattern recognition, while System 2 (Schema Engine) models HPC-PFC functions by extracting, storing, and applying schemas with priorities. (B) The information processing flow during inference, where the meta-trained transformer performs iterative decomposition of complex expressions (e.g., “walk right twice after turn left"), guided by the schema library. With each pass, System 1 processes one layer of the composition tree and stores bindings in episodic memory, while System 2 handles schema selection and expansion. This dual-process design enables Mirage to systematically solve compositional problems through rule-guided hierarchical processing.
2 Methodology
Intuition.
Mirage consists of two complementary components, an explicit, rule-based Schema Manager (System 2) and a meta-learned Transformer Decomposer (System 1), in a tight feedback loop that mirrors the HPC-PFC division of labor: rapid episodic binding versus deliberate schema-level manipulation. Figure 2 sketches the pipeline; below we formalize its elements and interactions.
2.1 Problem setting
We consider supervised sequence-to-sequence learning from an input space $\mathcal{X}$ to an output space $\mathcal{Y}$, where both are sequences of tokens. Each training example $(x,y)\!\in\!\mathcal{X}\!\times\!\mathcal{Y}$ is generated by a term rewriting system $\mathcal{G}$ (a “grammar”). Term rewriting systems consist of rewrite rules $\mathbf{x}\!\mapsto\!\mathbf{y}$ specifying that the string $\mathbf{x}$ rewrites into $\mathbf{y}$, where each string can contain variables. When such rules are allowed to apply cyclically to their own outputs, this formalism becomes Turing-complete, making it a powerful and generic starting point. We can fully specify such term rewriting system via a finite set of schemas. A schema is an $r$-ary function
$\sigma_{M}:(t_{1},\dots,t_{r})\;\longmapsto\;s_{1}\dots\ s_{q}\ M\,s_{q+1}\,\dots\,s_{r}$
where every placeholder or argument $t_{i}$ is bound at application time to a token or sub-schema $s_{j}$ in composition. A schema is triggered in a sequence by a corresponding modifier token, $M$. For example, in the SCAN grammar, $\sigma_{\textsc{after}}(a,b)=a\;\textit{after}\;b$, where after represents the modifier token. A schema may be composed of multiple sub-schemas, too, becoming a compound schema. For example, we might specify $\sigma_{\textsc{after-twice}}(a,b,c)=a\;\textit{after}\;b\;\textit{twice}$. A grammar may contain compound schemas, but must contain a standard atomic schema for every recognized modifier token.
To preserve deterministic compositional generation and resolvement, we must order the schemas in a grammar by tagging them with an integer priority $\pi(\sigma)$ (e.g. “turn left” must be executed before “twice”). This necessity may be best explained through the simple analogy to ‘PEMDAS’ order of operations in mathematics. Specific operations must take priority to ensure consistent solutions.
Lastly, a grammar must contain an evaluator which executes some function on its arguments. While a schema outputs a sequence, an evaluator $\psi$ maps a schema and arguments to an output sequence that may or may not consist of arguments to the schema.
$\psi:(\sigma,(t_{1},\dots,t_{r}))\;\longmapsto\;e_{1}\dots\ e_{n}$
Hence a grammar is the 3-tuple $\mathcal{G}=(\Sigma,\pi,\psi)$ with $\Sigma$ the schema set. Through recursion, a grammar can output sequences of arbitrary length.
2.2 System 1: Meta-Learned Transformer Decomposer
System 1 is a decoder-only Transformer $\mathcal{T}_{\theta}$ that does a single step of compositional reasoning, and which we apply cyclically to its own outputs to solve reasoning problems. Specifically, it is trained to perform single-step decomposition of any input sequence generated by an arbitrary grammar $\mathcal{G}$, given the definition of $\mathcal{G}$ in-context. This allows compositional problems to be reasoned about in sequential steps, with more complex (i.e. deeper) compositional problems requiring more steps. At inference time, this results in repeated applications of the System 1 transformer to fully decompose a compositional problem and return a result.
Breaking a compositional reasoning task into steps in this manner allows small models to execute complex compositional tasks without explicitly modeling compositional logic directly. Instead, models must only execute ‘prioritized pattern matching,’ or the identification of schema input patterns according to schema priority. This problem is more tractable than modeling k-deep composition, yet is sufficient to achieve systematic compositionality when used iteratively.
2.3 Training Algorithm: Meta-Learning Generalizable Single-Step Decomposition
Sequences with composition can be represented as composition trees. Every non-leaf node in a composition tree built from a sequence must refer to a modifier or action and every leaf node must refer to a primitive. Each SCAN sequence represents a single composition tree. Single-step decomposition for a compositional sequence is equivalent to reducing the depth-level of the composition tree by 1 after resolving all sub-trees that extend into that deepest level.
To parameterize a general function for single-step decomposition of any composition tree defined by a valid grammar, we meta-learn $\mathcal{T}_{\theta}$ on an endless stream of randomly generated grammars $\{\mathcal{G}_{m}\}_{m=1}^{\infty}$ (cf. Lake and Baroni (2018, 2023); Appendix A.1A.2). Training samples to the model consist of the ordered concatenation of the following 3 components: 1) A token sequence representation for a grammar as a set of schemas with corresponding priorities, 2) a maximum 2-deep composition input sequence of modifiers and atomic primitive tokens, generated by the grammar, and 3) the input sequence with single-step decomposition applied, denoted by replacing modifiers with schema names.
$\underbrace{\text{schema-library}}_{\texttt{Schema 1}\,\cdots\texttt{Schema M}}\;\texttt{LP\_SEP}\;\underbrace{\text{input sequence}}_{\smash{\textstyle x}}\;\texttt{SEP}\;\underbrace{\text{target sequence}}_{\smash{\text{DECOMPOSE}(x)}}\;\texttt{EOS}$
2.4 Inference Algorithm: Iterative Refinement with Replacement
Given a trained $\mathcal{T}_{\theta}$, we can reconstruct output sequences from input sequences of any composition depth and generated by any arbitrary grammar. In summary, we take input sequence $x^{(k)}$ with $k$ levels of schema composition and apply our model $\mathcal{T}_{\theta}$. Given the model output, we search for ‘schema name tokens’, which the transformer outputs to represent an instance of decomposition being applied. For every schema name token, we replace that token, as well as all of the arguments required for the schema referenced by that name, with a single atomic primitive. We store this replacement in a ‘replacement map’ (simple hash map or table). This results in a new sequence, $x^{(k-1)}$, with $k-1$ levels of composition (the last layer or bottom layer of the composition tree has been removed). We then apply the model to the new sequence after replacement and decompose the next layer. This zero-shot-given-a-grammar inference procedure is fully defined by Algorithm 1.
@@ 2.4 Inference Algorithm: Iterative Refinement with Replacement
6:  $y\leftarrow\mathcal{T}_{\theta}\bigl{(}x^{(d)}\bigr{)}$triangleright predict next-level decomposition
@@
9:   $p\leftarrow\texttt{ApplySchema}(\sigma_{s},\,a_{1},\dots,a_{m})$triangleright collapse into one primitive
@@
12:  $x^{(d-1)}\leftarrow y$triangleright one less level of composition
@@
After receiving the output from algorithm 1, we simply resolve or unwind the complete value for $S_{final}$, replacing any tokens in the sequence that match keys in $\lambda$ with their corresponding value in $\lambda$.
The number of model iterations required to completely resolve a schema is equal to the depth of the composition tree, where the root node sits at depth 0. Consider an input sequence of length $N$, with a max number of arguments per schema of $A$ (in SCAN, $A=2$). Then, the depth of the tree is $O(\log_{A}N)$, and as such, the number of decomposition or reasoning steps is generally logarithmic in sequence length. This allows processing extremely large compositional sequences in a small number of reasoning steps.
2.5 System 2 details: complementary schema extractors
Schema extraction is treated as a modular component whose internal strategy can be varied without altering the rest of Mirage. We currently provide two contrasting implementations: a symbolic search procedure and a learnable graph-based model built atop a recent computational model of the hippocampus George et al. (2021). Their different strategies make it unlikely that both will fail on the same input, and either one can be replaced as improved algorithms or neuroscientific insights become available.
Extractor Option One: CSCG-inspired extractor
Following the neuroscientific inspiration of the two-component model, we utilize Clone-Structured Causal Graph (CSCG) George et al. (2021), effective in modeling cognitive maps in mice Sun et al. (2025). Adapted from Cloned Hidden Markov Models (CHMM) Dedieu et al. (2019), CSCGs efficiently represent complex sequences via a discrete latent state model. CSCGs further constrain each latent state to deterministically emit a single token, effectively ‘cloning’ each token into a large number of hidden states. Critically, CSCGs can rebind these emission matrices to generalize to patterns analogous to those they were trained on, allowing them to learn simple algorithms Swaminathan et al. (2023). Our approach works by extending this rebinding capability.
The original rebinding mechanism triggers upon encountering unexpected tokens and depends on stable “anchor” tokens that are rarely rebound. We extend this mechanism to bidirectional contexts where novel primitives might precede anchors. Thus, we introduce an adapted rebinding algorithm to explicitly detect anchors and rebinding slots within such contexts, defined formally below.
[…]
1. Episode decoding. We train a CHMM on the raw token stream. Viterbi decoding yields a latent state trace for each demonstration, forming episodes $e=\langle\mathbf{x},\mathbf{y},\mathbf{z}\rangle$ where $\mathbf{z}$ is the state sequence. These episodes serve as structured surrogates of hippocampal encodings.
2. Schema discovery and pruning. Candidate schemas are generated by aligning episode pairs $(e_{i},e_{j})$ using the longest common subsequence (LCS) under a constraint enforcing consistent variable bindings. A candidate is accepted if it validates on all supporting episodes via forward simulation and is supported by at least min_support examples, filtering out ambiguous or redundant rules. Surviving schemas are deduplicated using a keyword–coverage heuristic and pruned iteratively if compositionally subsumed by higher-support schemas. We visualize the set of extracted schemas in Appendix Figure 4.

### v3

@@ Abstract
Deep learning models struggle with systematic compositional generalization, a hallmark of human cognition. We propose Mirage, a neuro-inspired dual-process model that offers a processing account for this ability. It combines a fast, intuitive “System 1” (a meta-trained Transformer) with a deliberate, rule-based “System 2” (a Schema Engine), mirroring the brain’s neocortical and hippocampal–prefrontal circuits. Trained to perform general, single-step decomposition on a stream of random grammars, Mirage achieves >>99% accuracy on all splits of the SCAN benchmark in a task-agnostic setting. Ablations confirm that the model’s systematic behavior emerges from the architectural interplay of its two systems, particularly its use of explicit, prioritized schemas and iterative refinement. In line with recent progress on recursive/recurrent Transformer approaches, Mirage preserves an iterative neural update while externalizing declarative control into an interpretable schema module. Our work provides a concrete computational model for interpreting how compositional reasoning can arise from a modular cognitive architecture.
@@ 1 Introduction
Humans effortlessly generalize by composing known concepts in new ways, a capacity termed systematic compositionality [1, 2]. Yet, even large-scale neural networks often fail at such generalization, tending to learn shallow heuristics instead of abstract rules [3, 4, 5]. This paper provides a processing account for compositional reasoning, inspired by the dual-process architecture of the human brain.
Cognitive neuroscience suggests that intelligent behavior arises from the interplay of complementary learning systems (CLS) [6, 7]. This theory posits that a fast, intuitive “System 1,” supported by the neocortex, handles pattern recognition, while a slower, deliberate “System 2,” involving the hippocampus (HPC) and prefrontal cortex (PFC), supports structured, episodic, and goal-directed reasoning [8, 9]. The HPC rapidly encodes specific experiences (episodes), which the PFC abstracts over time into reusable schemas, generalizable knowledge structures, that enable flexible planning and inference [10, 11]. This division of labor allows for both rapid adaptation and systematic application of knowledge.
We instantiate this theory in Mirage (Meta-Inference with Rules and Abstractions from Generalized Experience), a dual-process model with two interacting components:
@@
    System 1 (Neural Decomposer): A compact Transformer trained via meta-learning to perform a single step of compositional decomposition. This parallels the fast, pattern-matching function of the neocortex.
@@
    System 2 (Schema Engine): An explicit symbolic module that extracts, stores, and applies prioritized schemas, managing variable bindings in an episodic memory. This mirrors the deliberate, schema-based reasoning of the HPC–PFC loop.
At inference, these systems engage in an iterative loop: System 2 extracts a compositional grammar from examples and uses it to guide the repeated application of System 1. Each pass of the System 1 Transformer refines the problem representation by one level, allowing the model to recursively decompose complex structures. This architecture provides an interpretable mechanism for generalization. On the SCAN benchmark [4], Mirage achieves state-of-the-art performance across all splits, despite its Transformer component never being trained on SCAN-specific data. Our work demonstrates that a modular, neuro-inspired architecture can provide a powerful and interpretable processing account for systematic generalization.
Figure 1: Mirage architecture and inference loop. (A) System 1 (Neural Decomposer) is a meta-trained Transformer for pattern matching. System 2 (Schema Engine) extracts and manages a library of prioritized schemas, modeling HPC–PFC function. (B) During inference, the systems iterate. For a command like “walk right twice after turn left,” System 2 provides the relevant grammar. System 1 applies one decomposition step per pass, with outputs stored in an episodic memory. This process repeats, reducing the composition tree layer by layer until a primitive action sequence is produced.
2 A Dual-Process Framework
Mirage solves compositional tasks through an iterative refinement process managed by its two systems. The core idea is to offload declarative knowledge (the "what") to an explicit schema library (System 2) and procedural knowledge (the "how") to a general-purpose neural reasoner (System 1).
System 2: Schema Engine.
A compositional task is defined by a grammar $\mathcal{G}=(\Sigma,\pi,\psi)$, where $\Sigma$ is a set of schemas (rewrite rules), $\pi$ is a priority ordering, and $\psi$ is an evaluator for primitive actions. System 2’s role is to induce this grammar from a small set of examples. This extraction is modular; we show it can be done with either a symbolic rule-miner or a neuro-inspired model based on Clone-Structured Causal Graphs (CSCGs) [12, 13]. For details on these methods, see Section˜B.2 and Appendix˜E.
System 1: Meta-Learned Transformer Decomposer.
System 1 is a decoder-only Transformer $\mathcal{T}_{\theta}$ trained not to solve any single task, but to perform a generic, single step of compositional decomposition. We meta-learn $\mathcal{T}_{\theta}$ on an endless stream of randomly generated grammars. Each training sample consists of a grammar definition, an input sequence, and the target output after one decomposition step at the highest-priority location. This forces the model to learn the abstract *process* of rule application rather than memorizing solutions for a specific task. See Section˜A.3 and Section˜A.4 for a full description of the training procedure and model vocabulary.
Inference as a Processing Account.
At inference time, System 2 first extracts the task grammar $\mathcal{G}$. Then, for a given input command, the model enters a loop (Figure˜1B) that serves as a processing account of step-by-step deliberation. In each step, the current sequence and the grammar are fed to the Transformer $\mathcal{T}_{\theta}$. For example, given ‘walk right twice after turn left’ and a grammar where ‘twice’ has higher priority than ‘after’, the Transformer identifies ‘walk right twice’ and rewrites it to a placeholder token, say P1. The sequence becomes ‘P1 after turn left’. In the next iteration, it decomposes this simpler structure. This process repeats until only a sequence of primitive actions remains. This iterative refinement allows a small model with a fixed context window to solve problems of arbitrary compositional depth. The full inference algorithm is detailed in Algorithm˜1.
3 Experiments and Behavioral Analysis
We evaluate Mirage on the canonical SCAN benchmark [4], which tests systematic generalization to novel compositions and longer sequences. The System 1 Transformer (1.19M parameters) is trained task-agnostically on random grammars; for evaluation, the System 2 Schema Engine extracts the SCAN grammar from its training split, which is then provided to the frozen Transformer.
3.1 Compositional Generalization Performance
As shown in Table˜1, Mirage achieves near-perfect accuracy on all SCAN splits, including those designed to foil standard sequence-to-sequence models. A vanilla Transformer trained directly on SCAN’s simple split masters it but completely fails on the generalization splits, indicating it learns superficial correlations rather than abstract rules. Providing the schema library as a simple prompt (‘Transformer+SC_Library’) does not help, confirming that a specialized architecture and training objective are required to properly utilize schematic knowledge.
Table 1: SCAN accuracy (mean  $\pm$ SEM over 4 runs). Baselines are re-trained on each split’s training set. Mirage uses a single, task-agnostically trained Transformer for all splits.
SCAN splits
Model Full Task Simple Length Add prim. ‘jump’ Template
Transformer N/A 99.8599.85± $\pm\,0.00$ 13.5813.58± $\pm\,0.01$ 0.400.40± $\pm\,0.13$ 3.093.09± $\pm\,3.01$
Transformer+SC_Library N/A 99.9199.91± $\pm\,0.11$ 15.8615.86± $\pm\,1.36$ 0.030.03± $\pm\,0.04$ 0.000.00± $\pm\,0.00$
Mirage (ours) 99.5999.59± $\pm\,0.24$ 99.5099.50± $\pm\,0.30$ 99.3599.35± $\pm\,0.41$ 99.6599.65± $\pm\,0.20$ 99.5599.55± $\pm\,0.23$
3.2 Processing Account: Ablation Studies
To understand which components drive this behavior, we conducted several ablations. These studies confirm that Mirage’s performance is an emergent property of its cognitive architecture, not just one component.
  1. 1.
    Schema priorities are crucial for cognitive control. Removing the explicit priority ordering from the schema library causes accuracy to plummet from 99.6% to 71.9%. The model becomes unable to resolve ambiguous commands (e.g., ‘turn left twice and walk’), analogous to a failure of executive function in selecting the correct order of operations.
  2. 2.
    Iterative refinement enables deliberation. Replacing the iterative loop with a single-pass, end-to-end mapping fails to generalize. The step-by-step process is essential, allowing the model to focus on one sub-problem at a time, a form of deliberate, sequential reasoning that monolithic models lack.
  3. 3.
    Accurate schemas are necessary. Introducing minor errors into the extracted grammar (e.g., misidentifying a schema’s arguments) causes performance to collapse to nearly 0%. This highlights the model’s reliance on accurate declarative knowledge. Its reasoning algorithm is sound, but it cannot function with a corrupted world model, a behavior consistent with human reasoning [14].
4 Related Work
Research on compositionality spans auxiliary objectives, neuro-symbolic models, meta-learning, and analyses of Transformer mechanisms. Auxiliary supervision can encourage structural understanding in Transformers [15], while memory-augmented models like LANE [16] explore variable-slot reasoning. Neuro-symbolic systems such as the Compositional Program Generator [17] and the Neural-Symbolic Recursive Machine [18] highlight the benefits of grammar-based modularity, paralleling Mirage’s schema-based design.
Meta-learning approaches have also demonstrated strong potential for systematic generalization. In particular, Lake and Baroni [19] show that a meta-learned neural network can achieve human-like compositional generalization, supporting the view that inductive biases learned from experience can yield flexible reasoning skills.
Transformer-focused studies reveal both strengths and limits: while large models can sometimes achieve compositional generalization [20], their reasoning often emerges only through grokking [21] or limited mechanistic patterns [22], underscoring the need for specialized architectures.
Finally, neuroscience perspectives propose that hippocampal replay supports compositional computation [23], and that HPC–PFC circuits represent an evolutionary solution to flexible reasoning [24]. These ideas motivate Mirage’s dual-process design, bridging cognitive neuroscience and AI.
5 Discussion and Conclusion
Mirage offers a hybrid approach to compositional generalization, using a neural network for a general procedural task (decomposition) while relying on an explicit, interpretable symbolic layer for declarative knowledge and control. Unlike monolithic models that must implicitly learn reasoning algorithms through “grokking,” Mirage explicitly encodes a process of deliberate, iterative refinement. This dual-process framework offers a concrete processing account for systematic generalization. By instantiating cognitive theories of complementary learning systems, our model achieves state-of-the-art zero-shot performance on SCAN. We release the full implementation of Mirage on GitHub. Our main contributions as an interpretive work are:
  • Processing Account: We show that systematicity can emerge from an architecture that separates procedural skill from declarative knowledge and applies them iteratively.
  • Behavioral Account: The model’s success where standard Transformers fail demonstrates the behavioral advantage of a neuro-inspired, modular design.
  • Developmental Account: Meta-learning provides a hypothesis for how agents can acquire robust, generalizable cognitive algorithms from varied experience.
  • Interpretability Account: Because Mirage explicitly marks and resolves schemas step by step, it produces a transparent trace of intermediate problem states. This contrasts with black-box models that leap directly to an answer, making our framework inherently more interpretable and diagnostically useful.
Connection to Cognitive Learning Systems
Beyond enabling stepwise compositional inference, schema formation as utilized by Mirage provides a mechanism for rapid assimilation of new rules into structured knowledge. In neuroscience, once an associative schema is established, new information consistent with that schema is integrated and becomes neocortically supported unusually quickly [10, 11, 25]. Recent work further links medial prefrontal schema representations with reinforcement-learning style updating and deployment for flexible behavior [9]. Mirage’s prioritized schema selection echoes this PFC-mediated control over hippocampal content during retrieval and planning [11].
In addition, when viewed through a continual-learning lens, schemas could act as structured inductive priors: previously consolidated structure may be used to constrain and accelerate learning on new tasks. Bayesian treatments formalize this as using posteriors from past tasks to define priors for future ones, while also highlighting practical limits when approximations are poor [26, 27]. These observations motivate future work in the direction of probabilistic extension: inferring over candidate schema targets (and, when needed, over the library itself) to maintain sample-efficient generalization when direct schema identification is challenging.
Relation to recurrent/recursive Transformers
A growing line of work explores iterative neural reasoning with weight-tying over steps, truncated credit assignment, and, in some cases, adaptive halting, yielding recurrent or recursive Transformer loops [28, 29, 30, 31]. Mirage lies squarely within this trend by retaining an iterative per-step neural update, yet it differs in where the control signal lives: rather than learning control implicitly, we externalize it as explicit schemas with priorities. This keeps the empirical benefits of looped refinement while producing stepwise, editable traces and an algorithmic-level account of the reasoning process, which are central to Cognitive Interpretability. In this sense, Mirage complements recurrent and recursive Transformer practice with a cognitively motivated control layer, clarifying how and why each step is chosen.
Limitations and Future Work
While successful on SCAN, the current framework relies on reliable extraction of a consistent grammar. Future work should explore robustness to noisy or incomplete schemas. Such issues could potentially be addressed by extending the extractor with probabilistic rule learning and/or the Neural Decomposer with active retrieval mechanisms. We also plan to apply this architecture to more complex reasoning tasks and use its inspectable states to test specific neuroscientific hypotheses about HPC–PFC interactions during planning. By building explicit, interpretable cognitive architectures, we can move beyond black-box evaluations and toward a mechanistic understanding of reasoning in neural systems.
Appendix A Methodology details
A.1 Formal definition of Transformer Inference
@@
6:  $y\leftarrow\mathcal{T}_{\theta}\bigl(x^{(d)}\bigr)$ $\triangleright$ predict next-level decomposition
@@
9:   $p\leftarrow\texttt{ApplySchema}(\sigma_{s},\,a_{1},\dots,a_{m})$ $\triangleright$ collapse into one primitive
@@
12:  $x^{(d-1)}\leftarrow y$ $\triangleright$ one less level of composition
@@
A.2 Formal definition of CSCG-based System 2 Algorithm